# An Introduction to Zero-Order Optimization Techniques for Robotics

### v1

@@ II-B3 Gaussian Smoothing
Alternatively, one may sample from the surface of a unit sphere [21]. This technique guarantees that the norm of the direction vectors is bounded.
To summarize, although the assumptions in [18, 20, 12] vary, the underlying ideas are of the same nature; sampling along random directions may match the performance of gradient descent while requiring $n$ times less function evaluation.
II-C Simulated Annealing
So far, we have seen that performing random approximate gradient descent can be sample efficient in the convex setting. However, one may ask whether those algorithms can converge to global solutions in nonconvex settings, as it is the case for the Greedy Local Search introduced in Algorithm 2. It is well known that, in its vanilla form, gradient descent is prone to local minima. However, additive Gaussian noise can lead to asymptotic global convergence guarantees. To understand why, we need to study the following Langevin dynamics:
$dX_{t}$ $=\nabla_{x}\log p(X_{t})+dW_{t}$ (11)
Here $p$ is a probability distribution and $W_{t}$ is a Brownian motion. $\nabla_{x}\log p$ is the score function of the distribution. Note that the gradient is taken with respect to the input of $p$ (not with respect to parameters of $p$ as in the log-likelihood trick of Gaussian smoothing). This Stochastic Differential Equation (SDE) has been extensively studied and it can be shown that $X_{t}$ converges to the distribution defined by $p$ [22]. Consequently, this SDE can be used to sample from a distribution with probability density $p$. One way to leverage this result in the context of global optimization is to choose the following probability distribution:
$p\propto\exp\left(-\frac{1}{\lambda}f\right)$ (12)
$\nabla_{x}\log p(x)=-\frac{1}{\lambda}\nabla_{x}f(x)$ (13)
Therefore, this SDE can be discretized in order to derive a recursive algorithm. Following this idea, [24] shows that with an appropriate schedule of step sizes $\alpha_{k}$ and $\gamma_{k}$, the following discretized Stochastic Langevin Gradient Dynamics guarantees asymptotic convergence to global solutions.
$x_{k+1}=x_{k}-\alpha_{k}g_{k}+\gamma_{k}w_{k}$ (14)
where $w_{k}$ is Gaussian noise and where $g_{k}$ is of the form $(\nabla f(x_{k})+\xi_{k})$ with $\xi_{k}$ representing stochastic noise. Hence, $g_{k}$ can be interpreted as a noisy approximation of the gradient. Intuitively, the Gaussian noise $w_{k}$ helps to escape local minima. The step sizes $\alpha_{k}$ and $\gamma_{k}$ are chosen such that the corresponding SDE has a temperature $\lambda$ that tends slowly to zero. This is called Simulated Annealing. Intuitively, slowly decreasing the temperature will guide the iterates towards a global solution.
These results can be used to prove the global convergence of SPSA. More precisely, the iterative scheme in Equation (14) with $g$ as the SPSA estimate defined in Equation (4) converges to a global minimum [25]. In fact, in its default version (i.e. without additive noise), SPSA converges to a global minimum [25]. This is because the noise of the estimate (i.e. $\xi_{k}$) is enough to escape local minima and to achieve global convergence. We are not aware of such results in the context of randomized smoothing. However, we can conjecture that similar results may be valid.
To summarize, we have seen that stochasticity can help gradient descent algorithms to reach global solutions. Although these results are highly relevant theoretically, they have limited practical impact as they are only asymptotic.
Remark 3.
At first sight, it might be hard to relate the discretized Langevin update to the Greedy Local Search in Algorithm 2. However, following the observation from [26], the if statement can be related to the Metropolis-Hastings algorithm [23]. We refer the reader to the Appendix A for more details.
@@ III Trajectory Optimization
In this section, we show how random search allows us to understand derivative-free optimization algorithms commonly used in robotics. In practice, Trajectory Optimization aims to solve
$\min_{\{u_{0},u_{1},\dots u_{T-1}\}}$ $\;\sum_{t=0}^{T-1}c_{t}(x_{t},u_{t})+c_{T}(x_{T})$ (15)
such that $\,x_{0}=x,\,\mbox{and}\;\;x_{t+1}=f_{\text{dyn}}(x_{t},u_{t})$
Here $x$ denotes the state, $u$ the control variable, $(c_{t})_{t}$ the running cost functions, $c_{T}$ the terminal cost function, $T$ the time horizon and $f_{\text{dyn}}$ the dynamics. Note that we abuse the notation by using $x$ to denote the state instead of the optimization variable. Importantly, the constraints are implicit; the optimization is only done with respect to the control variables. Therefore, this problem is unconstrained and called the single shooting approach. In the end, it can be written in the form of Problem (1).
@@ III-A Predictive sampling
In robotics, Predictive sampling [9] is one of the most simple random search approach. This method can be seen as a variation of the Greedy Local Search (Algorithm 2). Instead of sampling one search direction, it samples $K$ search directions and keeps the best. If no direction provides an improvement, the previous estimate is kept. Algorithm 4 summarizes the procedure.
@@ 6      $x\leftarrow\operatorname{argmin}_{\tilde{x}\in D}f(\tilde{x})$
Importantly, all the function evaluations can be performed in parallel. While this approach can be quite effective, recent MPC demonstrations rely on more sophisticated algorithms that use gradient information.
@@ III-B The log-sum-exp transform: MPPI
Model Predictive Path Integral (MPPI) [10] is a derivative-free TO method initially derived from an information theoretic point of view. MPPI iteratively samples around the current guess and updates it using a weighted exponential average. More specifically, given a current guess $x$, MPPI samples $K$ variables $x_{k}$ following a Gaussian distribution, $\mathcal{N}(x,\Sigma)$. Then, it takes an update according to the following rule:
$x\leftarrow\sum_{k=1}^{K}w_{k}x_{k}$ (16)
@@
$w_{k}=\frac{\exp(-\frac{1}{\lambda}(f(x_{k})-\rho))}{\sum_{j=1}^{K}\exp(-\frac{1}{\lambda}(f(x_{j})-\rho))}$ (17)
@@ 3       Compute weights $w_{1},\dots w_{K}$
Let’s see how this update rule can be interpreted as an instance of Gaussian smoothing. Instead of the surrogate function defined in Equation 5, let’s consider the (continuous) log-sum-exp function.
$f_{\mu,\lambda}(x)=-\lambda\log\left(\mathbb{E}\left[\exp\left(-\frac{1}{\lambda}f(x+\mu\epsilon)\right)\right]\right)~{}$ (18)
where $\epsilon\sim\mathcal{N}(0,\Sigma)$ and where $\lambda>0$ is called the temperature. As $\mu$ goes to zero, we recover $f$. Therefore, the gradients of this surrogate can be used to approximate the gradients of the original function, $f$. Furthermore, as $\lambda\rightarrow\infty,f_{\mu,\lambda}\rightarrow f_{\mu}$ [15]. Consequently, this surrogate can be seen as a generalization of Gaussian smoothing. The gradient of the surrogate reads:
$\nabla f_{\mu,\lambda}(x)=\frac{-\lambda\mathbb{E}\left[\exp\left(-\frac{1}{\lambda}f(x+\mu\epsilon)\right)\Sigma^{-1}\epsilon\right]}{\mu\mathbb{E}\left[\exp\left(-\frac{1}{\lambda}f(x+\mu\epsilon)\right)\right]}~{}$ (19)
@@
$g=\frac{-\lambda}{\mu}\frac{\sum\limits_{k=1}^{K}\exp\left(-\frac{1}{\lambda}f(x+\mu\epsilon_{k})\right)\Sigma^{-1}\epsilon_{k}}{\sum\limits_{j=1}^{K}\exp\left(-\frac{1}{\lambda}f(x+\mu\epsilon_{j})\right)}$ (20)
Similarly to the randomized smoothing case, this update is invariant to additive translation of the function $f$. Let’s consider $\mu=1$. We will show that MPPI follows a natural gradient step [27]:
$x\leftarrow x-\alpha F^{-1}g$ (21)
where $\alpha$ is the step size and where $F$ is the Fisher information matrix which is equal to the inverse of the covariance matrix for Gaussian distributions. The idea of the natural gradient is to build an update step that is invariant to changes of variables. More details on the natural gradient are provided in the appendix B. If $\alpha=\frac{1}{\lambda}$, we have
@@
$=\sum\limits_{i=1}^{K}w_{k}(x+\epsilon_{k})$ (22)
As $(x+\epsilon_{k})\sim\mathcal{N}(x,\Sigma)$, we recover the MPPI update. Furthermore, if $\Sigma=\sigma^{2}I$, then $\frac{\sigma^{2}}{\lambda}\leq\frac{1}{L}$ where $L$ is the Lipschitz function of the surrogate function [15]. Therefore, the step size chosen by MPPI can be interpreted as a conservative estimate of the standard optimal step size from convex optimization [28].
Figure 1 illustrates how the log-sum-exp smoothing differs from Nesterov’s randomized smoothing. This simple example function has two types of local minima: one minima is global but narrow and the other one is local but wide. Interestingly, Gaussian smoothing gives similar importance to both minima; the two minima now have a similar depth and width. In contrast, the log-sum-exp smoothing gives more importance to the global minima; the global minima is much wider and almost absorbs the local one.
Figure 1: randomized smoothing (RS) compared to log-sum-exp (LSE) smoothing with $\lambda=0.1$.
To the best of our knowledge, [15] proposes the first explanation of the success of MPPI from an optimization perspective. [15] shows that the log-sum-exp smoothing can be interpreted as an interpolation between the default randomized smoothing and the Moreau envelope. The idea is that the Moreau envelope is a better approximation than RS but is harder to compute. MPPI elegantly trades-off approximation quality and computational complexity. At the cost of more samples, relying on this surrogate can improve convergence rates [15]. Intuitively, this estimator requires many samples because of the presence of the expectation in the denominator in Equation (19). With the progress of parallel computing, this requirement of multiple samples might not be a limitation at all. However, using many samples will render the update step essentially deterministic, which will make the algorithm prone to getting trapped in local minima.
@@
The log-sum-exp smoothing reminds of the risk-seeking control formulation [29]. We refer the reader to Appendix C for a discussion on the topic.
@@
The log-sum-exp smoothing reminds of the simulated annealing idea. Similarly to simulated annealing, as the temperature $\lambda$ goes to 00, the surrogate concentrates around the global minima. In fact,
$p\propto\exp\left(-\frac{1}{\lambda}f_{\mu,\lambda}\right)=\mathbb{E}\left[\exp\left(-\frac{1}{\lambda}f(x+\mu\epsilon)\right)\right]$
@@ III-C Covariance Matrix Adaptation (CMA)
So far, we have kept the covariance matrix $\Sigma$ fixed across various iterations. However, in practice, it is not necessarily clear how to choose this parameter according to the problem. In this section, we show how to derive a Covariance Matrix Adaptation (CMA) scheme. Interestingly, the optimization of the smooth surrogate function can be interpreted as an optimization in the space of Gaussian probability distribution. Indeed, optimizing Equation (5) can be interpreted as a search over the mean of a Gaussian distribution. More specifically, we can write
$f_{\mu}(x)=\mathbb{E}[f(x+\mu u)]=\mathbb{E}_{z\sim\mathcal{N}(x,\Sigma)}\left[f(z)\right]$ (23)
Therefore, Gaussian smoothing can be seen as the search of a Gaussian distribution that minimizes the expectation of $f$ under that distribution. This distribution can be interpreted as the belief of where the global minimum of the function $f$ might be. A natural generalization is to not only optimize the mean but also the covariance. More specifically, one can apply gradient descent to the following problem:
$\min_{\theta=(x,\Sigma)}J(\theta)$ (24)
@@
$J(\theta)=\mathbb{E}_{z\sim\mathcal{N}(x,\Sigma)}\left[f(z)\right]$ (25)
$J(\theta)$ is minimal when the distribution $\mathcal{N}(x,\Sigma)$ concentrates around the minima of $f$ [30]. Note that this idea do not necessarily need to rely on Gaussian distributions. Similarly to the randomized smoothing case, we can use the log-likelihood trick:
$\nabla J(\theta)=\mathbb{E}_{z\sim\mathcal{N}(x,\Sigma)}\left[f(z)\nabla_{\theta}\log(p_{\theta}(z))\right]$ (26)
Here, the gradient of the probability distribution is taken with respect to both the mean and the covariance. Akimoto et al. [31] shows that the natural gradient update $\Delta\theta=F(\theta)^{-1}\nabla J(\theta)$ can be written:
$\Delta\Sigma$ $=\mathbb{E}_{z\sim\mathcal{N}(x,\Sigma)}\left[f(z)(\left(z-x\right)\left(z-x\right)^{\top}-\Sigma)\right].$
$\Delta m$ $=\mathbb{E}_{z\sim\mathcal{N}(x,\Sigma)}\left[f(z)(z-x)\right]$ (27)
Similarly to the randomized smoothing case, this update step averages quantities that only rely on function evaluation of the original function. Following the natural gradient is especially relevant in this setting, as it is crucial that the update does not depend on the parametrization choice of the Gaussian distribution. We refer the reader to the appendix B for more details on the natural gradient. In practice, this update step can be approximated by samples $x_{k}\sim\mathcal{N}(x,\Sigma)$. Denoting $w_{k}=f(x_{k})$, the approximate update with a step size $\alpha$ reads
$\Sigma$ $\leftarrow(1-\alpha\sum\limits_{k=1}^{K}w_{k})\Sigma+\alpha\sum\limits_{k=1}^{K}w_{k}(x_{k}-x)(x_{k}-x)^{\top}$
$x$ $\leftarrow(1-\alpha\sum\limits_{k=1}^{K}w_{k})x+\alpha\sum\limits_{k=1}^{K}w_{k}x_{k}$ (28)
This allows us to recovers the CMA scheme (see Algorithm 6). Note that it is important to update the covariance matrix first as the covariance update must rely on the old $x$ (i.e. the one used to sample). As with Gaussian smoothing, a constant can be subtracted to $f$ without changing the expectation in Equation (26). Therefore, another choice of weights could be $w_{k}=f(x_{k})-f(x)$ as this renders the update invariant to translations of $f$. One could go further and try to design an algorithm invariant to any increasing transformation of the objective function. This can be done by replacing $f(x_{k})$ by ordered weights $w_{1}\leq w_{2}\leq\dots\leq w_{K}$. As pointed out by [31, 32, 30], this recovers the CMA-ES algorithm (without evolution path) [8]. [30] provides a theoretical justification for this choice of weights.

### v2

@@ II-C Simulated Annealing
At first glance, it might be hard to relate the discretized Langevin update to the Greedy Local Search in Algorithm 2. However, following the observation from [30], the if statement ensuring a strict decrease can be related to the Metropolis-Hastings algorithm [27]. We refer the reader to Appendix A for more details.
@@ III Trajectory Optimization
In this section, we show how random search allows us to understand derivative-free optimization algorithms that are commonly used in robotics. In practice, Trajectory Optimization aims to solve
$\min_{\{u_{0},u_{1},\dots u_{T-1}\}}$ $\;\sum_{t=0}^{T-1}c_{t}(x_{t},u_{t})+c_{T}(x_{T}),$ (17)
such that $\,x_{0}=x,\,\mbox{and}\;\;x_{t+1}=f_{\text{dyn}}(x_{t},u_{t}).$
Here $x$ denotes the state, $u$ the control variable, $(c_{t})_{t}$ the running cost functions, $c_{T}$ the terminal cost function, $T$ the time horizon, and $f_{\text{dyn}}$ the dynamics. Note that we abuse the notation by using $x$ to denote the state instead of the optimization variable. Importantly, the constraints are implicit; the optimization is performed only with respect to the control variables. Therefore, this problem is unconstrained and is referred to as the single shooting approach. Ultimately, it can be written in the form of Problem (1). The dimension of the search space is $n=Tn_{u}$, where $n_{u}$ denotes the dimension of the control inputs.
@@ III-A Predictive sampling
In robotics, Predictive sampling [9] is one of the simplest random search approaches. This method can be seen as a variation of the Greedy Local Search (Algorithm 2). Instead of sampling one search direction, it samples $K$ search directions and keeps the best. If no direction provides an improvement, the previous estimate is kept. Algorithm 4 summarizes the procedure.
@@ 5      Add $x+d_{k}$ to $D$
Importantly, all the function evaluations can be performed in parallel. While this approach can be quite effective, recent sampling-based MPC demonstrations typically rely on more sophisticated algorithms [2, 3].
@@ III-B The log-sum-exp transform: MPPI
Model Predictive Path Integral (MPPI) [10] is a derivative-free TO method initially derived from an information-theoretic perspective. MPPI iteratively samples around the current guess and updates it using a weighted exponential average. More specifically, given a current guess $x$, MPPI samples $K$ variables $x_{k}$ following independent Gaussian distributions centered in $x$ (i.e. $x_{k}\sim\mathcal{N}(x,\Sigma)$). Then, it performs an update according to the following rule:
$x\leftarrow\sum_{k=1}^{K}w_{k}x_{k},$ (18)
@@
$w_{k}=\frac{\exp(-\frac{1}{\lambda}(f(x_{k})-\rho))}{\sum_{j=1}^{K}\exp(-\frac{1}{\lambda}(f(x_{j})-\rho))},$ (19)
@@ 3   Compute weights $w_{1},\dots w_{K}$
Let’s see how this update rule can be interpreted as a variation of Gaussian smoothing. Instead of the surrogate function defined in Equation (5), let’s consider the (continuous) log-sum-exp transform function.
$f_{\mu,\lambda}(x)=-\lambda\log\left(\mathbb{E}\left[\exp\left(-\frac{1}{\lambda}f(x+\mu\epsilon)\right)\right]\right),$~ (20)
where $\epsilon\sim\mathcal{N}(0,\Sigma)$ and where $\lambda>0$ is a scalar called the temperature. As $\mu$ goes to zero, we recover $f$. Therefore, the gradients of this surrogate can be used to approximate the gradients of the original function, $f$. Furthermore, as $\lambda\rightarrow\infty,f_{\mu,\lambda}\rightarrow f_{\mu}$ [15]. Consequently, this surrogate can be seen as a generalization of Gaussian smoothing. The gradient of the surrogate reads:
$\hskip-2.84544pt\nabla f_{\mu,\lambda}(x)=\frac{-\lambda\mathbb{E}\left[\exp\left(-\frac{1}{\lambda}f(x+\mu\epsilon)\right)\Sigma^{-1}\epsilon\right]}{\mu\mathbb{E}\left[\exp\left(-\frac{1}{\lambda}f(x+\mu\epsilon)\right)\right]}.$~ (21)
@@
$g=\frac{-\lambda}{\mu}\frac{\sum\limits_{k=1}^{K}\exp\left(-\frac{1}{\lambda}f(x+\mu\epsilon_{k})\right)\Sigma^{-1}\epsilon_{k}}{\sum\limits_{j=1}^{K}\exp\left(-\frac{1}{\lambda}f(x+\mu\epsilon_{j})\right)}.$ (22)
Similarly to the randomized smoothing case, this update is invariant to additive translation of the function $f$. Let’s consider $\mu=1$. We show that MPPI follows a natural gradient step [31]:
$x\leftarrow x-\alpha F^{-1}g,$ (23)
where $\alpha$ is the step size, and where $F$ is the Fisher information matrix, which is equal to the inverse of the covariance matrix for Gaussian distributions. The idea of the natural gradient is to build an update step that is invariant to changes of variables. More details on the natural gradient are provided in Appendix B. If $\alpha=\frac{1}{\lambda}$, we have
@@
$=\sum\limits_{i=1}^{K}w_{k}(x+\epsilon_{k}).$ (24)
As $(x+\epsilon_{k})\sim\mathcal{N}(x,\Sigma)$, we recover the MPPI update. Furthermore, if $\Sigma=\sigma^{2}I$, then $\frac{\sigma^{2}}{\lambda}\leq\frac{1}{L}$, where $L$ is the Lipschitz function of the surrogate function [15]. Therefore, the step size chosen by MPPI can be interpreted as a conservative estimate of the standard optimal step size from convex optimization [32].
Figure 1 illustrates how the log-sum-exp smoothing differs from Nesterov’s randomized smoothing. The expectations are estimated with $100,000$ samples. This simple example function has two types of local minima: one minimum is global but narrow, and the other is non-global but wide. Interestingly, Gaussian smoothing gives similar importance to both minima; the two minima now have similar depths and widths. In contrast, the log-sum-exp smoothing gives more importance to the global minimum; the global minimum is much wider and almost absorbs the local one.
Figure 1: Randomized smoothing (RS) compared to log-sum-exp (LSE) smoothing (with $\Sigma=I$).
To the best of our knowledge, [15] propose the first explanation of the success of MPPI from an optimization perspective. [15] shows that the log-sum-exp smoothing can be interpreted as an interpolation between the default randomized smoothing and the Moreau envelope. The idea is that the Moreau envelope is a better approximation than RS but is harder to compute. MPPI elegantly trades off approximation quality and computational complexity. At the cost of more samples, relying on this surrogate can improve convergence rates [15]. Intuitively, this estimator requires many samples because of the presence of the expectation in the denominator in Equation (21). With the progress of parallel computing, this requirement for multiple samples might not be a limitation at all. However, using many samples renders the update step essentially deterministic, which makes the algorithm prone to getting trapped in local minima.
@@
The log-sum-exp smoothing resembles the risk-seeking control formulation [33]. We refer the reader to Appendix C for a discussion on the topic.
@@
The log-sum-exp smoothing is reminiscent of the simulated annealing idea. Similarly to simulated annealing, as the temperature $\lambda$ approaches 0, the surrogate concentrates around the global minima. In fact,
$p\propto\exp\left(-\frac{1}{\lambda}f_{\mu,\lambda}\right)=\mathbb{E}\left[\exp\left(-\frac{1}{\lambda}f(x+\mu\epsilon)\right)\right].$
@@ III-C Covariance Matrix Adaptation (CMA)
So far, we have kept the covariance matrix $\Sigma$ fixed across various iterations. However, in practice, it is not necessarily clear how to choose this parameter according to the problem. In this section, we show how to derive a Covariance Matrix Adaptation (CMA) scheme. Interestingly, the optimization of the smooth surrogate function can be interpreted as an optimization in the space of Gaussian probability distributions. Indeed, optimizing Equation (5) can be interpreted as a search over the mean of a Gaussian distribution. More specifically, we can write:
$f_{\mu}(x)=\mathbb{E}[f(x+\mu u)]=\mathbb{E}_{z\sim\mathcal{N}(x,\Sigma)}\left[f(z)\right].$ (25)
Therefore, Gaussian smoothing can be seen as the search for a Gaussian distribution that minimizes the expectation of $f$ under that distribution. This distribution can be interpreted as the belief of where the global minimum of the function $f$ might be. A natural generalization is to not only optimize the mean but also the covariance. More specifically, one can apply gradient descent to the following problem:
$\min_{\theta=(x,\Sigma)}J(\theta),$ (26)
@@
$J(\theta)=\mathbb{E}_{z\sim\mathcal{N}(x,\Sigma)}\left[f(z)\right].$ (27)
$J(\theta)$ is minimal when the distribution $\mathcal{N}(x,\Sigma)$ concentrates around the minima of $f$ [34]. Note that this idea does not necessarily need to rely on Gaussian distributions. Similarly to the randomized smoothing case, we can use the log-likelihood trick:
$\nabla J(\theta)=\mathbb{E}_{z\sim\mathcal{N}(x,\Sigma)}\left[f(z)\nabla_{\theta}\log(p_{\theta}(z))\right].$ (28)
Here, the gradient of the probability distribution is taken with respect to both the mean and the covariance. [35] shows that the natural gradient update, $\Delta\theta=F(\theta)^{-1}\nabla J(\theta)$}, can be written:
$\Delta\Sigma$ $=\mathbb{E}_{z\sim\mathcal{N}(x,\Sigma)}\left[f(z)(\left(z-x\right)\left(z-x\right)^{\top}-\Sigma)\right],$
$\Delta m$ $=\mathbb{E}_{z\sim\mathcal{N}(x,\Sigma)}\left[f(z)(z-x)\right].$ (29)
Similarly to the randomized smoothing case, this update step averages quantities that rely solely on function evaluations of the original function. Following the natural gradient is especially relevant in this setting, as it is crucial that the update does not depend on the choice of parameterization of the Gaussian distribution. We refer the reader to Appendix B for more details on the natural gradient. In practice, this update step can be approximated by samples $x_{k}\sim\mathcal{N}(x,\Sigma)$. Denoting $w_{k}=f(x_{k})$, the approximate update with a step size $\alpha$ reads
$\Sigma$ $\leftarrow(1-\alpha\sum\limits_{k=1}^{K}w_{k})\Sigma+\alpha\sum\limits_{k=1}^{K}w_{k}(x_{k}-x)(x_{k}-x)^{\top},$
$x$ $\leftarrow(1-\alpha\sum\limits_{k=1}^{K}w_{k})x+\alpha\sum\limits_{k=1}^{K}w_{k}x_{k}.$ (30)
This allows us to recover the CMA scheme (see Algorithm 6). Note that it is important to update the covariance matrix first, as the covariance update must rely on the old $x$ (i.e. the one used to sample). As with Gaussian smoothing, a constant can be subtracted from $f$ without changing the expectation in Equation (28). Therefore, another choice of weights could be $w_{k}=f(x_{k})-f(x)$, as this renders the update invariant to translations of $f$. One could go further and design an algorithm invariant to any increasing transformation of the objective by replacing $f(x_{k})$ with arbitrary weights $w_{k}$ sorted so that their order matches that of the function values $f(x_{k})$. As pointed out by [35, 36, 34], this recovers the CMA-ES algorithm (without evolution path) [8]. [34] provides a theoretical justification for this choice of weights.